3) by the action of vitamin D is formed from cholecalciferol (D3) by The active form of vitamin D is formed first and then th help plz asap
Answer #1

30) option(d) is correct.(liver,kidney)

Cholecalciferol is converted in the liver to Calcifediol.Calcifediol is further hydrolylated by kidney to form Calcitriol,the biological active form of vitamin D.

31)option(b) is correct.(3 pounds)

I consume 1 mocha latte coffee which gives 215 calories

So in 52 weeks ,I consume 52 mocha little coffee which gives 215× 52= 11,180 calories

3500 calories is equal to 1 pound of body fat.

So 11,180 calories= 11180÷3500= 3.19 pounds

32) option(c) is correct.(Beta carotene)

Eating too many carrot can bring too much beta carotene the molecules responsible for carrot's bright orange hue and a precusor of vitamin A .This can lead to excess carotene which can cause discolour skin.

33)option(b) is correct.(Bullimia nervosa.)

Bulimia nervosa is a serious potentially life treating eating disorders.People with bulimia may secretly binge- eating large amount of food with a lot of control over eating and then purge,trying to get rid of the extra calories in an unhealthy way.